OpenStudy (anonymous):

Which expression is the first step you take when finding the quotient of five over seven and three over two? five over seven times three over two five over seven times two over three seven over five times three over two seven over five times two over three

OpenStudy (anonymous):

The second, assuming this is: \(\LARGE\frac{\frac{5}{7}}{\frac{3}{2}}\), because you multiply by the inverse of the denominator.